The length of an American flag that is displayed at a government office is 3in less than twice its width. The area is 1710 square inches. Find the length and the width of the flag
Answer by jorel555(1290)

Let w be the width of the flag. Then the length would be 2w-3. So:
W (2w-3)=1710
2w^2-3w-1710=0
(2w+57)(w-30)=0
w=-57/2, 30
The width of the flag is 30 in.; the length is 57 in....😊😊😊😊